# Reborn: Reinforcement-Learned Boundary Segmentation with Iterative Training for Unsupervised ASR

### v1

@@ Abstract
Unsupervised automatic speech recognition (ASR) aims to learn the mapping between the speech signal and its corresponding textual transcription without the supervision of paired speech-text data. A word/phoneme in the speech signal is represented by a segment of speech signal with variable length and unknown boundary, and this segmental structure makes learning the mapping between speech and text challenging, especially without paired data. In this paper, we propose Reborn, Reinforcement-Learned Boundary Segmentation with Iterative Training for Unsupervised ASR. Reborn alternates between (1) training a segmentation model that predicts the boundaries of the segmental structures in speech signals and (2) training the phoneme prediction model, whose input is a segmental structure segmented by the segmentation model, to predict a phoneme transcription. Since supervised data for training the segmentation model is not available, we use reinforcement learning to train the segmentation model to favor segmentations that yield phoneme sequence predictions with a lower perplexity. We conduct extensive experiments and find that under the same setting, Reborn outperforms all prior unsupervised ASR models on LibriSpeech, TIMIT, and five non-English languages in Multilingual LibriSpeech. We comprehensively analyze why the boundaries learned by Reborn improve the unsupervised ASR performance. [^2] Our code will be available soon.
Machine Learning, ICML
@@ 1 Introduction
Automatic speech recognition (ASR) systems convert speech signals into their transcription texts. Most state-of-the-art (SoTA) ASR systems are supervisedly trained with a large amount of labeled speech-text paired data (Chen et al., 2021; Radford et al., 2023). However, more than 80% of languages in the world have limited access to speech-text paired data (Pratap et al., 2023), and collecting such paired data requires intensive labor and costs, limiting the applicability of supervised ASR systems to these languages. This leads to significant efforts devoted to developing unsupervised ASR (UASR) systems, which aim to learn the mapping between speech signals and textual transcriptions (words or phonemes) without any speech-text pairs (Liu et al., 2018; Chen et al., 2019; Baevski et al., 2021; Liu et al., 2022; Gao et al., 2023).
@@
UASR models learn to align the distribution of input speech signal and output text without paired data. Learning to match two distributions of sequences unsupervisedly has been extensively studied in unsupervised neural machine translation (NMT) (Artetxe et al., 2018), where the aim is to learn a neural network that can translate text in a source language to text in a target language without paired data (Lample et al., 2018a, b). Prior works show that adversarial training (Goodfellow et al., 2014) can be used to learn such mappings (Lample et al., 2018a, b), which employs a generator learning to translate and a discriminator learning to distinguish generated text from the real text in the target language.
@@
Can we adopt such an adversarial training scheme for UASR? In unsupervised NMT, text can be easily tokenized into sub-word tokens, so unsupervised NMT only needs to learn the mapping between the source and target language’s token embeddings. However, in UASR, a phoneme or word is represented by a variable-length segment in the speech signal whose boundaries are unknown. Moreover, the length of a speech signal is much longer than the length of its textual transcription. The above characteristics of speech make learning the mapping between the segmental structures in speech and textual transcription challenging. Existing works in unsupervised ASR rely on heuristics or handcrafted rules to obtain the boundaries of the segmental structures (Liu et al., 2018; Yeh et al., 2018; Baevski et al., 2021). Yet, such handcrafted boundaries are often sub-optimal, bottlenecking the performance of UASR.
@@
This paper focuses on obtaining better segmental boundaries to improve the mapping between segmental structure and textual transcription. We propose Reborn (Reinforcement-Learned Boundary Segmentation with Iterative Training), an unsupervised ASR framework with a segmentation model and a phoneme prediction model. The segmentation model determines segmental structure boundaries in speech signals, while the phoneme prediction model assigns a phoneme to each segmental structure. We use an iterative algorithm that alternates between two stages to train the segmentation model and the phoneme prediction model: The first stage trains the segmentation model. Since we do not have the ground truth segmental boundaries, we use reinforcement learning (RL) to provide unsupervised guidance for training the segmentation model. We experiment with various learning objectives and implement them as reward functions for learning the segmentation model. In the second stage, based on the boundaries predicted by the segmentation model learned in the previous stage, we use an adversarial loss similar to generative adversarial networks (GANs) (Goodfellow et al., 2014) to train the phoneme prediction model.
@@
We conduct extensive experiments on LibriSpeech (Panayotov et al., 2015), TIMIT (Garofolo, 1993), and Multilingual LibriSpeech (MLS) (Pratap et al., 2020) to compare the phoneme error rate (PER) and word error rate (WER) with prior works. We show that our method outperforms all prior UASR methods on LibriSpeech, TIMIT, and five languages in MLS when using the same amount of training data. We perform thorough ablation studies to show that iterative training and the rewards we design are critical to the performance of Reborn. By analyzing the segmental structure obtained by our segmentation model, we find that the segmental structures are acoustic units smaller than phonemes, which helps the phoneme prediction model predict more accurate transcriptions.
2 Related Work: Unsupervised ASR (UASR)
@@
An ASR model takes speech signals as the input and predicts the textual transcriptions. Unsupervised ASR aims to train an ASR model without access to paired speech-text data. Instead, the only available training data is unlabeled speech data and unlabeled text data, while the correspondence between speech and text is not available. In this paper, we follow prior UASR works (Liu et al., 2018; Baevski et al., 2021) to predict phoneme transcriptions from speech signals. Learning the mapping between speech signals and phoneme sequences can be formulated as a distribution matching problem, where we want to learn an ASR model whose output phoneme sequences match the distribution of real phoneme sequences.
@@
Yeh et al. (2018) solve this distribution matching problem by an unsupervised loss function based on the empirical output distribution matching (ODM) (Liu et al., 2017), which guides the ASR model to output phoneme sequences that have statistic distributions close to real phoneme sequences.
@@
Liu et al. (2018) propose to use the generative adversarial network (GAN) (Goodfellow et al., 2014) for UASR. GAN solves the distribution matching problem by training a generator whose output distribution resembles a target distribution, and a discriminator is trained to distinguish the outputs of the generator and the samples from the target distribution. When using GAN to train unsupervised ASR, the generator is the phoneme prediction model. The phoneme prediction model takes in speech features of the segmental structures in the speech signal, and outputs a phoneme transcription. Prior works rely on heuristics or separately trained unsupervised phoneme segmentation models to find the boundary of the segmental structures (Wang et al., 2018; Liu et al., 2018; Chen et al., 2019).
With the emergence of foundation models, wav2vec-U (Baevski et al., 2021) shows that UASR performance can greatly benefit from using speech foundation models as the feature extractor. However, wav2vec-U still relies on some heuristic feature preprocessing steps based on $k$-means to find the boundaries of the segmental structure in the speech signal. In our paper, we call the segmentation boundaries obtained by the feature preprocessing in wav2vec-U as ”$k$-means segmentation”. EURO (Gao et al., 2023) changes the decoding algorithm used to decode the phoneme prediction sequences and explores other speech foundation models for feature extraction, including HuBERT (Hsu et al., 2021) and WavLM (Chen et al., 2022).
In these prior methods, the segmentation is either non-trainable or learned separately from the phoneme prediction model. Reborn differs from prior works by introducing a trainable segmentation model tailored for the phoneme prediction model, and the segmentation model and phoneme prediction model can be iteratively learned. wav2vec-U 2.0 (Liu et al., 2022) simplifies the feature preprocessing step in wav2vec-U and does not explicitly consider the segmentation structure in their model. Reborn has performance better than wav2vec-U 2.0 on almost all datasets, showing that learning the segmentation boundary for feature preprocessing is important to the performance of UASR.
@@
Figure 1: (a) and (b): Reborn iterates between using RL to train the segmentation model and using adversarial training to train the phoneme prediction model. (c): An illustration of the segmentation post-processing. Given the original segmentation and the predicted phoneme sequence, we merge the segments that result in the same phoneme prediction into the same segment, yielding the post-processed boundary.
@@ 3 Method: Reborn
The difficulty of mapping speech signals to their corresponding phoneme transcriptions lies in the segmental structures in speech whose boundaries are unknown. To tackle this challenge, we propose Reborn, which trains a UASR system with unpaired speech and text data. Reborn contains a segmentation model and a phoneme prediction model. The segmentation model takes the speech feature as input and determines the boundaries of the segmental structures in the speech signal, and the phoneme prediction model predicts a phoneme for each segmental structure. We will use the terms segment and segmental structure interchangeably.
The overall training process of Reborn is outlined in Algorithm 1 and Figure 1. First, we initialize the phoneme prediction model using wav2vec-U (Section 3.3). Next, the training of Reborn iterates through two stages: Stage 1 (Figure 1(a), Section 3.1): Train the segmentation model to learn a better segmentation until the segmentation model converges while fixing the phoneme prediction model. Stage 2 (Figure 1(b), Section 3.2): Training the phoneme predictor based on the segment boundaries predicted by the segmentation model until the phoneme prediction model converges. The iterative process ceases when the UASR performance does not improve over the previous iteration. In UASR, we cannot use PER as a validation metric to select the best model or determine when to stop training. We use an unsupervised evaluation metric to select the best model, detailed in Appendix C.2.
@@ 3.1.1   Segmentation Model
Given an input speech signal, we use a self-supervised feature extractor model to extract the speech features $X=[x_{1},\cdots,x_{T}]$ from the waveform. The segmentation model takes the speech features and predicts the boundary of the segmental structure in the speech features. For each feature $x_{i}$, the segmentation model assigns a binary value $\hat{b}_{i}$, indicating whether $x_{i}$ is the first frame of a segment. We split $X$ into segments based on $\hat{b}_{1:T}$ and mean-pool the features within the same segment. The mean-pooled features $[s_{1},\cdots,s_{T^{\prime}}]$ are forwarded through the phoneme prediction model to obtain the phoneme prediction $[y_{1},\cdots,y_{T^{\prime}}]$. The resulting phoneme prediction will be de-duplicated into $[y^{\prime}_{1},\cdots,y^{\prime}_{M}]$ by removing the same consecutive phoneme.
@@ 3.1.2   Training the Segmentation Model with RL
We train the segmentation model using the policy gradient method (Sutton et al., 1999). For each utterance, we calculate an utterance-wise reward $R$ (defined in Eq. 4 in the next subsection) from the de-duplicated phoneme prediction to train the segmentation model. Based on the policy gradient method, the segmentation model $\pi_{\theta}$ is optimized using the following gradient: $\mathbb{E}\left[\sum_{t=1}^{T}\nabla_{\theta}\log\pi_{\theta}(\hat{b}_{t}|X)R\right]$, where $\mathbb{E}$ is taken over the training data distribution and approximated with the mean of a batch of training sequences.
@@ 3.1.3   Reward Designs
The perplexity difference reward is designed to favor phoneme segmentation better than the segmentation learned in the previous iteration. Intuitively, a better phoneme segmentation prediction $\hat{b}_{1:T^{\prime}}$ should yield a more reasonable phoneme prediction $y_{1:T^{\prime}}$. We use perplexity (PPL), the negative likelihood of a phoneme sequence scored by a phoneme language model (LM), to evaluate how reasonable a phoneme sequence is. Perplexity measures how likely a phoneme sequence can be observed in the real world; a phoneme sequence with a lower perplexity means it is more likely to be observed in real-world phoneme datasets. We use KenLM (Heafield, 2011) as the phoneme LM, where the LM is trained with the unlabeled text corpora. To guide the segmentation model to generate a segmentation with a better phoneme prediction than the phoneme predictions obtained in the previous iteration, we define perplexity difference reward as follows:
@@
where $\text{PPL}_{\theta-1}$ is the perplexity of $Y^{\prime}_{\theta-1}$ and $\text{PPL}_{\theta}$ is the perplexity of $Y^{\prime}_{\theta}$. The perplexity difference reward guides the segmentation model to produce segmentation that results in a phoneme sequence with lower perplexity.
When only using $R_{\text{ppl}}$, the segmentation model learns some segmentation method that leads to phoneme predictions with more $<$silence$>$ token, which has a lower perplexity but does not correspond to better phoneme prediction results. To avoid such unwanted behaviors, we design two regularization rewards, edit distance reward and length difference reward, to ensure that the policy learned by the segmentation model does not deviate too much from the segmentation learned in the previous iteration. The two regularization rewards punish the segmentation model if $Y^{\prime}_{\theta}$ is too different from $Y^{\prime}_{\theta-1}$. Crucially, regularization works only if the $Y^{\prime}_{\theta-1}$ in the first iteration is good enough. As previously mentioned, we ensure this by using the predictions from a trained wav2vec-U as $Y^{\prime}_{\theta-1}$ in the first iteration.
@@
We use Levenshtein distance $d_{\text{Lev}}$ as the edit distance to calculate the difference between $Y^{\prime}_{\theta-1}$ and $Y^{\prime}_{\theta}$. We take the normalized negative edit distance as the reward, which is defined as follows:
$R_{\text{edit}}=-\frac{d_{\text{Lev}}(Y^{\prime}_{\theta-1},Y^{\prime}_{\theta})}{|Y^{\prime}_{\theta-1}|}$ (2)
@@
We use the length difference reward $R_{\text{len}}$ to guide the segmentation model to predict segmentation such that the length of $Y^{\prime}_{\theta}$ does not differ significantly from the length of $Y^{\prime}_{\theta-1}$. The length difference reward for an utterance is defined as:
$R_{\text{len}}=1-\frac{\left||Y^{\prime}_{\theta}|-|Y^{\prime}_{\theta-1}|\right|}{|Y^{\prime}_{\theta-1}|},$ (3)
where $|Y^{\prime}|$ is the length of $Y^{\prime}$.
Utterance-Wise Reward
@@
where $c_{\text{ppl}}$, $c_{\text{edit}}$ and $c_{\text{len}}$ are the weighting coefficients. During training, $R_{\text{ppl}}$, $R_{\text{edit}}$ and $R_{\text{len}}$ are normalized within each training batch.Appendix C.2 discusses how the coefficients are determined without speech-text paired data.

### v2

@@ Abstract
Unsupervised automatic speech recognition (ASR) aims to learn the mapping between the speech signal and its corresponding textual transcription without the supervision of paired speech-text data. A word/phoneme in the speech signal is represented by a segment of speech signal with variable length and unknown boundary, and this segmental structure makes learning the mapping between speech and text challenging, especially without paired data. In this paper, we propose Reborn, Reinforcement-Learned Boundary Segmentation with Iterative Training for Unsupervised ASR. Reborn alternates between (1) training a segmentation model that predicts the boundaries of the segmental structures in speech signals and (2) training the phoneme prediction model, whose input is the speech feature segmented by the segmentation model, to predict a phoneme transcription. Since supervised data for training the segmentation model is not available, we use reinforcement learning to train the segmentation model to favor segmentations that yield phoneme sequence predictions with a lower perplexity. We conduct extensive experiments and find that under the same setting, Reborn outperforms all prior unsupervised ASR models on LibriSpeech, TIMIT, and five non-English languages in Multilingual LibriSpeech. We comprehensively analyze why the boundaries learned by Reborn improve the unsupervised ASR performance.
@@ 1 Introduction
Automatic speech recognition (ASR) systems convert speech signals into their transcription texts. Most state-of-the-art (SoTA) ASR systems are supervisedly trained with a large amount of labeled speech-text paired data [9, 49]. However, more than 80% of languages in the world have limited access to speech-text paired data [48], and collecting such paired data requires intensive labor and costs, limiting the applicability of supervised ASR systems to these languages. This leads to significant efforts devoted to developing unsupervised ASR (UASR) systems, which aim to learn the mapping between speech signals and textual transcriptions (words or phonemes) without any speech-text pairs [31, 10, 4, 28, 17].
@@
UASR models learn to align the distribution of input speech signal and output text without paired data. Learning to match two distributions of sequences unsupervisedly has been extensively studied in unsupervised neural machine translation (NMT) [1], where the aim is to learn a neural network that can translate text in a source language to text in a target language without paired data [24, 25]. Prior works show that adversarial training [19] can be used to learn such mappings [24, 25], which employs a generator learning to translate and a discriminator learning to distinguish generated text from the real text in the target language.
@@
Can we adopt such an adversarial training scheme for UASR? In unsupervised NMT, text can be easily tokenized into sub-word tokens, so unsupervised NMT only needs to learn the mapping between the source and target language’s token embeddings. However, in UASR, a phoneme or word is represented by a variable-length segment in the speech signal whose boundaries are unknown. Moreover, the length of a speech signal is much longer than the length of its textual transcription. The above characteristics of speech make learning the mapping between the segmental structures in speech and textual transcription challenging. Existing works in unsupervised ASR rely on handcrafted rules or separately learned modules to obtain the boundaries of the segmental structures [31, 60, 4]. Yet, such handcrafted boundaries are often sub-optimal, bottlenecking the performance of UASR.
@@
This paper focuses on learning better segmental boundaries to improve the mapping between segmented speech and textual transcription. We propose Reborn (Reinforcement-Learned Boundary Segmentation with Iterative Training), an unsupervised ASR framework with a segmentation model and a phoneme prediction model. The segmentation model determines segmental structure boundaries in speech signals, while the phoneme prediction model assigns a phoneme to each segmental structure. After properly initializing the phoneme prediction model, we use an iterative algorithm that alternates between two stages to train the segmentation model and refine the phoneme prediction model: The first stage trains the segmentation model. Since we do not have the ground truth segmental boundaries, we use reinforcement learning to train the segmentation model to favor segmentations that yield phoneme sequence predictions with a lower perplexity. We experiment with various learning objectives and implement them as reward functions for learning the segmentation model. In the second stage, based on the boundaries predicted by the segmentation model learned in the previous stage, we use an adversarial loss similar to generative adversarial networks (GANs) [19] to train the phoneme prediction model.
@@
We conduct extensive experiments on LibriSpeech [42], TIMIT [18], and Multilingual LibriSpeech (MLS) [47] to compare the phoneme/phone error rate (PER) and word error rate (WER) with prior works. We show that our method outperforms all prior UASR methods on LibriSpeech, TIMIT, and five languages in MLS when using the same amount of training data. We perform thorough ablation studies to show that iterative training and the rewards we design are critical to the performance of Reborn. By analyzing the segmental structure obtained by our segmentation model, we find that the segmental structures are acoustic units smaller than phonemes, which helps the phoneme prediction model predict more accurate transcriptions.
2 Related Work: Unsupervised ASR
@@
An ASR model takes speech signals as the input and predicts the textual transcriptions. Unsupervised ASR (UASR) aims to train an ASR model without access to paired speech-text data. Instead, the only available training data is unlabeled speech data and unlabeled text data, while the correspondence between speech and text is not available. In this paper, we follow prior UASR works [31, 4] to predict phoneme transcriptions from speech signals. Learning the mapping between speech signals and phoneme sequences can be formulated as a distribution matching problem, where we want to learn an ASR model whose output phoneme sequences match the distribution of real phoneme sequences.
@@
Yeh et al. [60] solve this distribution matching problem by an unsupervised loss function based on the empirical output distribution matching (ODM) [32], which guides the ASR model to output phoneme sequences that have statistic distributions close to real phoneme sequences.
@@
Liu et al. [31] propose to use the generative adversarial network (GAN) [19] for UASR. GAN solves the distribution matching problem by training a generator whose output distribution resembles a target distribution, and a discriminator is trained to distinguish the outputs of the generator and the samples from the target distribution. When using GAN to train unsupervised ASR, the generator is the phoneme prediction model. The phoneme prediction model takes in speech features of the segmental structures in the speech signal, and outputs a phoneme transcription. Prior works rely on hand-crafted or separately trained unsupervised phoneme segmentation models to find the boundary of the segmental structures [57, 31, 10].
@@
Figure 1: (a) and (b): Reborn iterates between using RL to train the segmentation model and using adversarial training to train the phoneme prediction model. (c): An illustration of the segmentation/boundary merging. Given the original segmentation and the predicted phoneme sequence, we merge the segments that result in the same phoneme prediction into the same segment, yielding the merged boundary.
With the emergence of foundation models, wav2vec-U [4] shows that UASR performance can greatly benefit from using speech foundation models as the feature extractor. However, wav2vec-U still relies on the two-stage feature preprocessing step based on $k$-means to find the boundaries of the segmental structure in the speech signal. In our paper, we call the segmentation boundaries obtained by the feature preprocessing in wav2vec-U as "$k$-means-based segmentation". EURO [17] changes the decoding algorithm used to decode the phoneme prediction sequences and explores other speech foundation models for feature extraction, including HuBERT [22] and WavLM [11].
In these prior methods, the segmentation is either non-trainable or learned separately from the phoneme prediction model. Reborn differs from prior works by introducing a trainable segmentation model tailored for the phoneme prediction model, and the segmentation model and phoneme prediction model can be iteratively polished. wav2vec-U 2.0 [28] simplifies the feature preprocessing step in wav2vec-U and does not explicitly consider the segmentation structure in their model. Reborn has performance better than wav2vec-U 2.0 on almost all datasets, showing that learning the segmentation boundary for feature preprocessing is important to the performance of UASR.
Although originated to be a cross-modality distribution matching problem, recent UASR is also related to representation learning. For example, the speech features might be extracted from self-supervised speech foundation models; and the segmentation problem is also commonly investigated in tasks like acoustic unit discovery. We leave the discussions about the two related topics in Appendix F.
@@ 3 Method: Reborn
The difficulty of mapping speech signals to their corresponding phoneme transcriptions lies in the segmental structures in speech whose boundaries are unknown. To tackle this challenge, we propose Reborn, which trains a UASR system with unpaired speech and text data. Reborn contains a segmentation model and a phoneme prediction model. The segmentation model takes the speech feature as input and determines the boundaries of the segmental structures in the speech signal, and the phoneme prediction model predicts a phoneme for each segmental structure. We will use the terms segment and segmental structure interchangeably. In our paper, we do not use the term segment in the exact sense as in linguistics, where segment refers to discrete units that can be identified in the stream of speech [14]. We use segment to broadly refer to a span in the speech, which may be a meaningful unit (e.g., a phone or word) or a span in the speech identified by the segmentation model (Section 3.1).
The overall training process of Reborn is outlined in Figure 1. First, we initialize the phoneme prediction model using wav2vec-U (Section 3.3). Next, the training of Reborn iterates through two stages: Stage 1 (Figure 1(a), Section 3.1): Training the segmentation model to learn a better segmentation until the segmentation model converges while fixing the phoneme prediction model. Stage 2 (Figure 1(b), Section 3.2): Training the phoneme predictor based on the segment boundaries predicted by the segmentation model until the phoneme prediction model converges. The iterative process ceases when the UASR performance does not improve over the previous iteration. In UASR, we cannot use PER as a validation metric to select the best model or determine when to stop training. We use an unsupervised evaluation metric to select the best model, detailed in Appendix E.4.
@@ 3.1.1 Segmentation Model
Given an input speech signal, we use a self-supervised feature extractor model to extract the speech features $X=[x_{1},\cdots,x_{T}]$ from the waveform. The segmentation model takes the speech features and predicts the boundary of the segmental structure in the speech features. For each feature $x_{i}$, the segmentation model assigns a binary value $\hat{b}_{i}$, indicating whether $x_{i}$ is the first frame of a segment. We split $X$ into segments based on $\hat{b}=\hat{b}_{1:T}$ and mean-pool the features within the same segment. The mean-pooled features $[s_{1},\cdots,s_{T^{\prime}}]$ are forwarded through the phoneme prediction model to obtain the phoneme prediction $[y_{1},\cdots,y_{T^{\prime}}]$. The resulting phoneme prediction will be de-duplicated into $[y^{\prime}_{1},\cdots,y^{\prime}_{M}]$ by removing the same consecutive phoneme.
@@ 3.1.2 Training the Segmentation Model with RL
We train the segmentation model using the policy gradient method [54] based on REINFORCE algorithm [59]. For each utterance, we calculate an utterance-wise reward $R$ (defined in Eq. 4 in the next subsection) from the de-duplicated phoneme prediction to train the segmentation model. Based on the policy gradient method, the segmentation model $\pi_{\theta}$ is optimized using the following gradient: $\mathbb{E}_{\hat{b}\sim\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}(\hat{b}|X)R\right]$, where $\mathbb{E}$ is taken over $\hat{b}$ sampled from $\pi_{\theta}$ and approximated with the mean of a batch of training sequences.
@@ 3.1.3 Reward Designs
The perplexity difference reward is designed to favor phoneme segmentation better than the segmentation learned in the previous iteration. Intuitively, a better phoneme segmentation prediction $\hat{b}_{1:T}$ should yield a more reasonable phoneme prediction $y^{\prime}_{1:M}$. We use perplexity (PPL), the negative likelihood of a phoneme sequence scored by a phoneme language model (LM), to evaluate how reasonable a phoneme sequence is. Perplexity measures how likely a phoneme sequence can be observed in the real world; a phoneme sequence with a lower perplexity means it is more likely to be observed in real-world phoneme datasets. We use a 4-gram phoneme LM trained on the phonemicized unlabeled text corpora. To guide the segmentation model to generate a segmentation with a better phoneme prediction than the phoneme predictions obtained in the previous iteration, we define perplexity difference reward as follows:
@@
where $\text{PPL}_{\theta-1}$ is the perplexity of $Y^{\prime}_{\theta-1}$ and $\text{PPL}_{\theta}$ is the perplexity of $Y^{\prime}_{\theta}$. The perplexity difference reward guides the segmentation model to produce segmentation that results in a phoneme sequence with lower perplexity comparing with the previous iteration.
When only using $R_{\text{ppl}}$, the segmentation model learns some segmentation method that leads to phoneme predictions with lower perplexity but does not correspond to better phoneme prediction results. To prevent such undesirable behaviors, we design two regularization rewards, the edit distance reward and length difference reward, to ensure that the policy learned by the segmentation model does not drastically alter the phoneme predictions between iterations. The two regularization rewards punish the segmentation model if $Y^{\prime}_{\theta}$ is too different from $Y^{\prime}_{\theta-1}$. Crucially, regularization works only if the $Y^{\prime}_{\theta-1}$ in the first iteration is good enough. As previously mentioned, we ensure this by using the predictions from a trained wav2vec-U as $Y^{\prime}_{\theta-1}$ in the first iteration.
@@
We use Levenshtein distance $d_{\text{Lev}}$ as the edit distance to calculate the difference between $Y^{\prime}_{\theta-1}$ and $Y^{\prime}_{\theta}$. We take the normalized negative edit distance as the reward, which is defined in Eq. 2
@@
We use the length difference reward $R_{\text{len}}$ to guide the segmentation model to predict segmentation such that the length of $Y^{\prime}_{\theta}$ does not differ significantly from the length of $Y^{\prime}_{\theta-1}$. The length difference reward for an utterance is defined in Eq. 3, where $|Y^{\prime}|$ is the length of $Y^{\prime}$.
$R_{\text{edit}}=-\frac{d_{\text{Lev}}(Y^{\prime}_{\theta-1},Y^{\prime}_{\theta})}{|Y^{\prime}_{\theta-1}|},$ (2)
$R_{\text{len}}=1-\frac{\left||Y^{\prime}_{\theta}|-|Y^{\prime}_{\theta-1}|\right|}{|Y^{\prime}_{\theta-1}|}$ (3)
@@
where $c_{\text{ppl}}$, $c_{\text{edit}}$ and $c_{\text{len}}$ are the weighting coefficients. During training, $R_{\text{ppl}}$, $R_{\text{edit}}$ and $R_{\text{len}}$ are normalized within each training batch. Appendix E.4 discusses how the coefficients are determined without speech-text paired data.